OpenStudy (anonymous):

what are the solutions of |x+6|>5 write the solutions as either the union or the intersection of two sets

OpenStudy (anonymous):

can someone help me with this please

OpenStudy (anonymous):

x+6=5 x+6=-5 now solve for x

OpenStudy (anonymous):

and because its greater sign it will be an OR

OpenStudy (anonymous):

the greater sign have the little line under it

OpenStudy (anonymous):

its still the same will still be an or

OpenStudy (anonymous):

x>-11

OpenStudy (anonymous):

I add 6and 5 am I doing it right

OpenStudy (anonymous):

yup so x=-11 and x=-1

OpenStudy (anonymous):

if u plug this into your original absolute value u will find these to be true

OpenStudy (anonymous):

:) does this make sense

OpenStudy (anonymous):

{x|x<-11}.....{x|>-1} is that how I write it

OpenStudy (anonymous):

their is the union cup facing down

OpenStudy (anonymous):

yes union

OpenStudy (anonymous):

union is facing up as like a U

OpenStudy (anonymous):

ok thnks

OpenStudy (anonymous):

\[(-\infty,-11) U (-1, \infty)\]

OpenStudy (anonymous):

thank you